91 adults selected randomly from a town,61 have health insurance. Find the 90% confidence interval for the true proportion of all adults in the town who have onsurance.
Answer by stanbon(75887)

91 adults selected randomly from a town,61 have health insurance. Find the 90% confidence interval for the true proportion of all adults in the town who have insurance.
----
sample proportion = p-hat = 61/91 = 0.67
------
ME = z*sqrt(p*q/n) = 1.645*sqrt(0.67*0.33/91) = 1.645*0.05 = 0.08
------------------------
90% CI:: 0.67-0.08 < p < 0.67+0.08
90% CI:: 0.59 < p < 0.75
-----------
Cheers,
Stan H.
-------------------